# Moduli spaces of untwisted wild Riemann surfaces

### v1

@@ 3.1. Definition and basic properties
Given integers $g\geq 0,m\geq 1$ and $\underline{\mathbf{d}}=(d_{\alpha,i})_{1\leq i\leq m,\alpha\in\Phi}$ with $d_{\alpha,i}\geq 0$, set $p=\max(\underline{\mathbf{d}})$. The wild mapping class group $\Gamma_{m,g}^{\mathfrak{t},\underline{\mathbf{d}}}$ is
$\Gamma_{m,g}^{\mathfrak{t},\underline{\mathbf{d}}}=\pi_{1}(\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p,\underline{\mathbf{d}}},x)$
for any point $x$.
3.1.2 Remark.
Since the stacks $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p,\underline{\mathbf{d}}}$ are path connected (because the same is true for $\mathsf{M}_{g,m}$ and for the fibres of the morphism $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p,\underline{\mathbf{d}}}\rightarrow\mathsf{M}_{g,m}$), the group $\Gamma_{m,g}^{\underline{\mathbf{d}}}$ is well defined up to isomorphism.
3.1.3. An explicit presentation of $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p,\underline{\mathbf{d}}}$
Recall that we can write $\mathsf{M}_{g,m}=[T_{g,m}/\Gamma_{g,m}]$, where the Teichmüller space $T_{g,m}$ is a contractible complex manifold (see section 4.1.9). By Theorem 2.4.6 the fibre product $\widetilde{WT}_{g,m}^{\mathfrak{t},\leq p,\underline{\mathbf{d}}}=\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p,\underline{\mathbf{d}}}\times_{\mathsf{M}_{g,m}}T_{g,m}$ is a complex manifold, and it is a locally trivial fibration over $T_{g,m}$ whose fibres are products of spaces of irregular types of root order $\mathbf{d}_{i}$ at a point, denoted by $\mathscr{I\!\!T}^{\mathfrak{t},\leq p,\mathbf{d}_{i}}$, for $1\leq i\leq m$. In fact, since $T_{g,m}$ is contractible the fibration $\widetilde{WT}_{g,m}^{\mathfrak{t},\leq p,\underline{\mathbf{d}}}\rightarrow T_{g,m}$ is topologically trivial: there is a homeomorphism
$T_{g,m}\times\prod_{i=1}^{m}\mathscr{I\!\!T}^{\mathfrak{t},\leq p,\mathbf{d}_{i}}\simeq\widetilde{WT}_{g,m}^{\mathfrak{t},\leq p,\underline{\mathbf{d}}}.$
The top row of the following cartesian diagram provides a chart of $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p,\underline{\mathbf{d}}}$:
[…]
3.1.4. Local versus global wild mapping class group
Let us denote the fundamental group of $\prod_{i=1}^{m}\mathscr{I\!\!T}^{\mathfrak{t},\leq p,\mathbf{d}_{i}}$ the “local” wild mapping class group - by $\mathrm{L}\Gamma^{\underline{\mathbf{d}}}$. The homotopy exact sequence attached to the right vertical map above yields a short exact sequence
$1\rightarrow\mathrm{L}\Gamma^{\underline{\mathbf{d}}}\rightarrow\Gamma_{m,g}^{\mathfrak{t},\underline{\mathbf{d}}}\rightarrow\Gamma_{g,m}\rightarrow 1$
expressing the wild mapping class group as an extension of the usual mapping class group by the local wild mapping class group.
3.2. Example: $\mathsf{WM}_{1,1}^{\mathfrak{t},\leq p}$
We have $\mathsf{M}_{1,1}=[\mathbf{H}/\mathrm{SL}_{2}(\mathbf{Z})]$, where $\mathbf{H}=\{\tau\in\mathbf{C}\mid\mathrm{Im}(\tau)>0\}$ and $\mathrm{SL}_{2}(\mathbf{Z})$ acts on it via Möbius transformations. Over $\mathbf{H}$ we have an elliptic curve (i.e. family of compact connected pointed Riemann surfaces of genus one)
$\mathcal{E}=(\mathbf{H}\times\mathbf{C})/\mathbf{Z}^{2},\;\sigma\colon\mathbf{H}\rightarrow\mathcal{E}$
where $(a,b)\in\mathbf{Z}^{2}$ acts sending $(\tau,z)$ to $(\tau,z+a\tau+b)$, and $\sigma(\tau)=(\tau,0)$. Therefore, the fibre $\mathcal{E}_{\tau}$ of $\mathcal{E}$ over $\tau\in\mathbf{H}$ is the torus $\mathbf{C}/(\mathbf{Z}+\tau\mathbf{Z})$ with the marked point 00. The coordinate $z$ is a local coordinate around each $\sigma(\tau)$, hence irregular types for $\mathcal{E}$ at $\sigma$ (with pole order at most $p$) can be written as
(3.2.0.1) $\sum_{j=1}^{p}z^{-j}A_{j}(\tau)$
where each $A_{i}\colon\mathbf{H}\rightarrow\mathfrak{t}$ is a holomorphic function. Therefore, we see that $\widetilde{WT}_{1,1}^{\mathfrak{t},\leq p}$ is isomorphic to $\mathbf{H}\times\mathfrak{t}^{p}$.
Division by $c\tau+d$ sends $\mathbf{Z}+\tau\mathbf{Z}=(c\tau+d)\mathbf{Z}+(a\tau+b)\mathbf{Z}$ to $\mathbf{Z}+\frac{a\tau+b}{c\tau+d}\mathbf{Z}$. If $\frac{a\tau+b}{c\tau+d}=\tau$, then $z\mapsto\frac{z}{c\tau+d}$ induces an automorphism of $\mathcal{E}_{\tau}$. Its action on irregular types can be explicitly written down using (3.2.0.1). For instance, take $\tau=i$ and $\begin{pmatrix}a&b\\ c&d\end{pmatrix}=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$. Then $\sum_{j=1}^{p}z^{-j}A_{j}(\tau)$ is sent to $\sum_{j=1}^{p}i^{j}z^{-j}A_{j}(\tau)$; in particular we see that, if $A_{j}(i)$ is non-zero for some $j$ not divisible by $4$, then the previous automorphism of $\mathcal{E}_{i}$ does not induce an automorphism of the relevant wild Riemann surface.
Generalising the above argument, one sees that points of $\mathsf{WM}_{1,1}^{\mathfrak{t},\leq p}$ have finitely many automorphisms (as the same is true for the elliptic curves $\mathcal{E}_{\tau}$), and that their automorphism groups are all trivial if for each $\tau$ there is $j$ not divisible by $2$ or $3$ such that $A_{j}(\tau)$ is non-zero.
4. Appendix: background on analytic stacks
@@
In this appendix we recall and give references for some generalities on stacks and their fundamental group which are needed in the text. The material in this section is not new, and is included for the reader’s convenience; for the same reason we try to provide some motivation for the concept of (analytic) stack. Needless to say, this is not meant to be an introduction to the subject, for which we refer the reader to the references given below.
4.1. Stacks
@@
We collect here the background material on stacks that is used in the text. We will focus on analytic stacks, and our main references will be [Noo05, BN06, Hei05].
4.1.1. Motivation: the moduli space of sections of a locally free sheaf
[…]
@@ 4.1.1. Motivation: the moduli space of sections of a locally free sheaf
4.1.2. Complex manifolds as sheaves
In section 4.1.1, we have described the vector bundle $V(\mathcal{E})$ attached to a locally free sheaf $\mathcal{E}$ in terms of the holomorphic maps from an arbitrary manifold $B$ to it: holomorphic maps $B\rightarrow V(\mathcal{E})$ are in natural bijection with couples $(\varphi,s)$ consisting of a holomorphic map $\varphi\colon B\rightarrow X$ and an element $s\in\Gamma(B,\varphi^{*}\mathcal{E})$. In general, for every complex manifold $X$, we can look at the functor $h(X)\colon\mathrm{Man}_{\mathbf{C}}^{op}\rightarrow\mathrm{Sets}$ sending $B$ to the set $\mathrm{Hom}(B,X)$ of holomorphic maps from $B$ to $X$. A holomorphic map $X\rightarrow Y$ induces a natural transformation of functors $h(X)\rightarrow h(Y)$. This gives rise to the Yoneda embedding

### v2

@@ 5.1. Wild algebraic curves
A wild algebraic curve of genus $g$ over a scheme $S$ with $m$ marked points, and pole orders bounded by $\underline{p}=(p_{1},\ldots,p_{m})$, is a triple $(\pi\colon C\rightarrow S,\underline{a},\underline{Q})$ consisting of a curve $\pi\colon C\rightarrow S$ of genus $g$, an $m$-tuple $\underline{a}=(a_{1},\ldots,a_{m})$ of mutually non-intersecting sections $a_{i}\colon S\rightarrow C$ of $\pi$, and an $m$-tuple $\underline{Q}=(Q_{1},\ldots,Q_{m})$ of irregular types $Q_{i}\in\Gamma(C,\mathfrak{t}\otimes_{\mathbf{C}}\mathscr{T}_{C,a_{i}}^{\leq p_{i}})$.
We denote by $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}(S)$ the groupoid of wild algebraic curves over $S$ of genus $g$ and pole orders bounded by $\underline{p}$ (morphisms being isomorphisms of $S$-schemes respecting the sections and the irregular types).
We obtain a 2-functor $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}\colon\mathrm{Sch}_{\mathbf{C}}\rightarrow\mathrm{Groupoids}$ sending $S$ to $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}(S)$, and a map $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}\rightarrow\mathcal{M}_{g,m}$ forgetting irregular types, where the target is the stack of $m$-pointed algebraic curves of genus $g$. The argument in the proof of Theorem 2.4.6(2) shows that this map is representable: if $S\rightarrow\mathcal{M}_{g,m}$ corresponds to a pointed curve $(C,\underline{a})$ then $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}\times_{\mathcal{M}_{g,m}}S=\prod_{i=1}^{m}\mathscr{I\!\!T}^{\mathfrak{t},\leq p_{i}}_{S,a_{i}}$, where the product is fibred over $S$.
5.1.2 Example.
Let $\mathfrak{g}=\mathfrak{sl}_{2}$ and $\underline{p}=(1,1,\ldots,1)$. For every $(\pi:C\rightarrow S,\underline{a})$, we have $\mathscr{T}_{C,a_{i}}^{\leq 1}=\mathcal{I}_{a_{i}}^{-1}/\mathcal{O}_{C}$. Seen as a sheaf on $S$, this is isomorphic to the inverse of $a_{i}^{*}(\Omega^{1}_{C/S})$ by [Sta18, Tag 06BB]. Therefore, in this case the line bundles of irregular types at each marked point generate a free subgroup of $\mathrm{Pic}(\mathcal{M}_{g,m})$ for $g\geq 3$; cf. [AC87, Theorem 2], where the line bundles $a_{i}^{*}(\Omega^{1}_{C/S})$ are denoted by $\psi_{i}$.
5.1.3 Lemma.
$\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}$ is an algebraic stack for the fppffppf topology.
Proof.
The properties in the statement hold true for $\mathcal{M}_{g,m}$, thanks to [LMB00, Proposition 4.5] and [Ols16, Example 4.4.13, Theorem 8.4.5] (resp. [Ols16, Theorem 13.1.2]) for $g\neq 1$ (resp. $g=1$; recall that $m\geq 1$ in this text). The fact that $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}$ is a stack follows as in the proof of Proposition 2.4.5, using [Sta18, Tag 023T]. Finally, as the map $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}\rightarrow\mathcal{M}_{g,m}$ is representable, by [LMB00, Proposition 4.5] the stack $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}$ is algebraic. ∎
5.1.4. Admissible wild algebraic curves
We stratify $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}$ according to the order of pole of the irregular types after evaluation at roots $\alpha\in\Phi$. Let $\underline{\mathbf{d}}=(d_{\alpha,i})_{\alpha\in\Phi,1\leq i\leq m}$ with $0\leq d_{\alpha,i}\leq p_{i}$. We define $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\leq\underline{\mathbf{d}}}$ to be the substack of $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}$ whose $S$-points are wild algebraic curves $(\pi\colon C\rightarrow S,\underline{a},\underline{Q})$ such that, for $1\leq i\leq m$, we have $\alpha\circ Q_{i}\in\Gamma(C,\mathscr{T}_{C,a_{i}}^{\leq d_{\alpha,i}})$. The map $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\leq\underline{\mathbf{d}}}\rightarrow\mathcal{M}_{g,m}$ is representable: the fibre product with an $S$-point of $\mathcal{M}_{g,m}$ corresponding to $(C,\underline{a})$ is the product over $S$ of the intersections of the schemes $\mathscr{I\!\!T}^{\mathfrak{t},\leq p_{i}}_{S,a_{i}}\times_{\mathscr{I\!\!T}^{\leq p_{i}}_{S,a_{i}}}\mathscr{I\!\!T}^{\leq d_{\alpha,i}}_{S,a_{i}}$, where $\mathscr{I\!\!T}^{\leq d_{\alpha,i}}_{S,a_{i}}$ denotes the scheme obtained setting $\mathbf{t}=\mathbf{C}$ at the beginning of section 5.1. Each $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\leq\underline{\mathbf{d}}}$ is a closed substack of $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}$; we define $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}\subset\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\leq\underline{\mathbf{d}}}$ to be the complement of the union of the substacks $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\leq\underline{\mathbf{d}^{\prime}}}$ for $\underline{\mathbf{d}^{\prime}}<\underline{\mathbf{d}}$. Its $S$-points are wild algebraic curves $(\pi\colon C\rightarrow S,\underline{a},\underline{Q})$ such that $Q_{i}$ has pole order at most $p_{i}$, each $\alpha(Q_{i})$ belongs to $\Gamma(C,\mathscr{T}_{C,a_{i}}^{\leq d_{\alpha,i}})$, and it does not belong to $\Gamma(C,\mathscr{T}_{C,a_{i}}^{\leq d_{\alpha,i}-1})$ anywhere on $S$ (i.e., the pullback of $\alpha(Q_{i})$ to every fibre $C_{s},s\in S$ does not belong to $\Gamma(C_{s},\mathscr{T}_{C_{s},a_{i,s}}^{\leq d_{\alpha,i}-1})$).
The fibre of the map $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}\rightarrow\mathcal{M}_{g,m}$ over a $\mathbf{C}$-point is an $m$-fold product of products of hyperplane complements as in Theorem 2.4.6(5). If this space is non-empty - a condition which only depends on $\underline{\mathbf{d}}$, and not on $\underline{p}$ - we say that $\underline{\mathbf{d}}$ is relevant.
5.1.5 Proposition.
Fix integers $g\geq 0,m\geq 1$, and tuples
$\underline{p}=(p_{1},\ldots,p_{m})\in\mathbf{Z}_{\geq 0}^{m},\quad\underline{\mathbf{d}}=(d_{\alpha,i})_{\alpha\in\Phi,1\leq i\leq m},\qquad 0\leq d_{\alpha,i}\leq p_{i},$
such that $\underline{\mathbf{d}}$ is relevant. The stack $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}$ is Deligne–Mumford if and only if
$2g-2+m+\sum_{i}\max(\mathbf{d}_{i})>0.$
Proof.
If $g\geq 2$ (resp. $g=1$) then the statement follows from [LMB00, Proposition 4.5] and the fact that $\mathcal{M}_{g}$ (resp. $\mathcal{M}_{1,1}$) is Deligne–Mumford by [Ols16, Theorem 8.4.5] (resp. [Ols16, Theorem 13.1.2]). If $g=0$ and $m\geq 3$ then $\mathcal{M}_{0,m}$ and $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}$ are algebraic varieties. Suppose that $g=0$ and $m=1$. As in section 4.2.3 we see that $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}$ is a quotient of a space of irregular types by the affine group. If $\max(\mathbf{d}_{1})\leq 1$ then $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}$ has complex points with automorphism group $\mathbf{G}_{a}$, therefore it is not Deligne–Mumford. If $\max(\mathbf{d}_{1})>1$ then the argument in section 4.2.3 shows that $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}$ is a locally closed substack of a weighted projective space, hence it is Deligne–Mumford. If $m=2$ and $\underline{\mathbf{d}}$ is identically zero then $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}$ has a substack isomorphic to $\mathcal{M}_{0,2}$ (parametrizing curves with trivial irregular types). Finally, suppose that $\underline{\mathbf{d}}=(\mathbf{d}_{1},\mathbf{d}_{2})$ is not identically 0. Assume, without loss of generality, that $\mathbf{d}_{2}\neq 0$; as in section 4.2.4 we have a representable map $\mathcal{WM}_{0,2}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}\rightarrow\mathcal{WM}_{0,2}^{\mathfrak{t},\leq\underline{p}^{\prime},\underline{\mathbf{d}}^{\prime}}$, hence it suffices to prove that $\mathcal{WM}_{0,2}^{\mathfrak{t},\leq\underline{p}^{\prime},\underline{\mathbf{d}}^{\prime}}$ is Deligne–Mumford. This is true because $\mathcal{WM}_{0,2}^{\mathfrak{t},\leq\underline{p}^{\prime},\underline{\mathbf{d}}^{\prime}}$ can be written, for the same reason as in section 4.2.4, as a quotient of a space of (non-zero) irregular types by a $\mathbf{G}_{m}$-action, making it a locally closed substack of a weighted projective stack. ∎
5.1.6. A parametrization of non-empty strata
One can parametrise the non-empty spaces $\mathcal{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}$ in terms of suitable filtrations of $\Phi$. To do so, we may take a fixed Riemann surface of genus $g$ with one marked point $a$, and look at a space $\mathscr{I\!\!T}^{\mathfrak{t},\leq p}$ of irregular types with pole at $a$ of order bounded above by a given integer $p\geq 0$; we want to parametrise non-empty strata of the root stratification on $\mathscr{I\!\!T}^{\mathfrak{t},\leq p}$.
Define a Levi subsystem of $\Phi$ to be a subset $\Phi^{\prime}\subset\Phi$ such that $\mathrm{span}_{\mathbf{C}}(\Phi^{\prime})\cap\Phi=\Phi^{\prime}$. A sequence $(\Phi_{1},\Phi_{2},\ldots)$ of Levi subsystems of $\Phi$ such that $\Phi_{i}\subset\Phi_{i+1}$ for $i\geq 1$ will be called a Levi filtration of $\Phi$. Such a filtration is of depth $d$ if $\Phi_{d}\subsetneq\Phi$ and $\Phi_{d+1}=\Phi$. We denote by $\mathcal{L}_{\Phi}^{\leq p}$ the set of Levi filtrations of $\Phi$ of depth at most $p$.
Let us identify $\mathscr{I\!\!T}^{\mathfrak{t},\leq p}$ with $\mathfrak{t}^{p}$, sending an irregular type $Q=\frac{A_{1}}{z}+\ldots+\frac{A_{p}}{z^{p}}$ to $(A_{1},\ldots,A_{p})$. We attach to $Q$ a Levi filtration obtained setting $\Phi_{i}=\{\alpha\in\Phi\mid\alpha(A_{j})=0\text{ if }i\leq j\leq p\}$ for $i\leq p$, and $\Phi_{i}=\Phi$ for $i>p$; in other words, the pole order of $\alpha(Q)$ equals $i\geq 0$ if and only if $\alpha$ belongs to $\Phi_{i+1}\smallsetminus\Phi_{i}$ (where we set $\Phi_{0}=\varnothing$). In particular, we see that the filtration $(\Phi_{i})_{i\geq 1}$ only depends on the root stratum to which $Q$ belongs.
Conversely, given a Levi filtration of depth at most $p$ we define $\mathbf{d}=(d_{\alpha})_{\alpha\in\Phi}$ setting $d_{\alpha}=i$ if $\alpha\in\Phi_{i+1}\smallsetminus\Phi_{i}$. To show that the stratum $\mathscr{I\!\!T}^{\mathfrak{t},\leq p,\mathbf{d}}$ is non-empty note that, if $\Phi_{i}\subsetneq\Phi_{i+1}$, then for each $\alpha\in\Phi_{i+1}\smallsetminus\Phi_{i}$ the intersection of the kernels of the roots in $\Phi_{i}$ meets the complement of $\ker(\alpha)$; therefore
$(\bigcap_{\alpha\in\Phi_{i}}\ker(\alpha))\cap(\bigcap_{\alpha\in\Phi_{i+1}\smallsetminus\Phi_{i}}(\mathfrak{t}\smallsetminus\ker(\alpha)))$
is non-empty (cf. [CFRW, Lemma 2.2.1]). Hence, there are $A_{1},A_{2},\ldots,A_{p}$ such that $\frac{A_{1}}{z}+\ldots+\frac{A_{p}}{z^{p}}$ belongs to $\mathscr{I\!\!T}^{\mathfrak{t},\leq p,\mathbf{d}}$.
To sum up, non-empty strata of the root stratification of $\mathscr{I\!\!T}^{\mathfrak{t},\leq p}$ are in bijection with $\mathcal{L}_{\Phi}^{\leq p}$.
Appendix A Background on analytic stacks
@@
In this appendix we recall and give references for some generalities on stacks and their fundamental group which are needed in the text. The material in this section is standard, and is included for the reader’s convenience; for the same reason we try to provide some motivation for the concept of (analytic) stack. Needless to say, this is not meant to be an introduction to the subject, for which we refer the reader to the references given below.
A.1. Stacks
@@
We collect here the background material on stacks that is used in the text. We will focus on analytic stacks, and our main references will be [Noo, BN06, Hei05].
A.1.1. Motivation: the moduli space of sections of a locally free sheaf
Let us start by showing the classical theory of moduli spaces in action from a modern viewpoint, in a simple example of interest in this document.
Fix a complex manifold $X$ and a locally free sheaf $\mathcal{E}$ of $\mathcal{O}_{X}$-modules of finite rank $n$. For every complex manifold $B$ with a map $\varphi\colon B\rightarrow X$ we can consider the global sections $\Gamma(B,\varphi^{*}\mathcal{E})$ of the pullback of $\mathcal{E}$ to $B$. We would like to construct the moduli space of all such global sections, i.e. a complex manifold $V(\mathcal{E})\xrightarrow{\pi}X$ such that there are natural bijections, for every $\varphi\colon B\rightarrow X$,
@@ A.1.1. Motivation: the moduli space of sections of a locally free sheaf
A.1.2. Complex manifolds as sheaves
In section A.1.1, we have described the vector bundle $V(\mathcal{E})$ attached to a locally free sheaf $\mathcal{E}$ in terms of the holomorphic maps from an arbitrary manifold $B$ to it: holomorphic maps $B\rightarrow V(\mathcal{E})$ are in natural bijection with couples $(\varphi,s)$ consisting of a holomorphic map $\varphi\colon B\rightarrow X$ and an element $s\in\Gamma(B,\varphi^{*}\mathcal{E})$.
In general, for every complex manifold $X$, we can look at the functor $h(X)\colon\mathrm{Man}_{\mathbf{C}}^{op}\rightarrow\mathrm{Sets}$ sending $B$ to the set $\mathrm{Hom}(B,X)$ of holomorphic maps from $B$ to $X$. A holomorphic map $X\rightarrow Y$ induces a natural transformation of functors $h(X)\rightarrow h(Y)$. This gives rise to the Yoneda embedding